Solving Problems Using Linear Systems
3. The girls' soccer team held a fundraising car wash. They charged $5 for each car and $8 for each van. They washed 44 cars and vans and collected $262. How many of ach type of vehicle did they wash?
(answers are: 30 cars and 14 vans, but i dont know how to solve it?)

Answer by stanbon(75887)

The girls' soccer team held a fundraising car wash. They charged $5 for each car and $8 for each van. They washed 44 cars and vans and collected $262. How many of ach type of vehicle did they wash?
------------------------------
Quantity Equation:
Quantity Equation: c + v = 44
Value EQuation...:5c +8v = 262
---------
Multiply thru the 1st equation by 5:
5c + 5v = 220
5c + 8v = 262
---------------
Subtract the 1st equation from the 2nd and solve for "v":
3v = 42
v = 14 (# of vans)
----------------------
Substitute into c+v = 44 to solve for "c":
c + 14 = 44
c = 30 (# of cars)
============================
Cheers,
Stan H.